OpenStudy (anonymous):

Find f(0) and f(2) for the given exponential function. f(x) = (1/4)^-x f(0)= f(2)=

OpenStudy (anonymous):

f(x)=(1/4)^-x f(0)=(1/4)^-0 = 1 anything to the 0 is 1 f(2) = (1/4)^-2 = 1/ (1/4)^2 = 1/(1/16) = 16